OpenStudy (anonymous):

Which equation could be used to solve the problem? Troy typed 375 word in 15 minutes. How many words per minute did he type (w)? A. w = 375 words • 15 minutes B. w = 375 words + 15 minutes C. w = 15 minutes ÷ 375 words D. w = 375 words ÷ 15 minutes

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

The key word here is 'per'. Think about a car going 95 miles per hour. We would write that as 95miles/1hour. Or maybe you ate 1 slice of pizza per 15 minutes. So we would write that as 1slice/15minutes. 'per' means to divide!

OpenStudy (mathmale):

This is a "rate" problem. Some rates that are probably familiar to you include $2.59/pound, two miles per minute, $1.25 each. You are given a strong hint in " How many words per minute did he type?" That word "per" requires you to divide. Divide what by what? Think about this.

OpenStudy (anonymous):

15

OpenStudy (anonymous):

Also remember that 375words/15minutes is a fraction. So you could simplify it and make the denominator (minutes) equal to 1. 375words/15minutes = 25words/1minute and that would be your true answer for 'words per minute'. But in this question they do not require you to simplify it, just choose which equation your would use

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

What is your reasoning behind choosing B?

OpenStudy (anonymous):

Note that how B is adding two different units 375 words + 15 minutes we can't just combine those two

OpenStudy (anonymous):

NO D.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which equation could be used to solve the problem? Oscar averaged 4 yards per carry in the football game. If he carried the ball 16 times, how many yards did he gain (y)? A. y = 4 • 16 B. y = 4 + 16 C. y = 4 ÷ 16 D. y = 16 ÷ 4

OpenStudy (anonymous):

@mathmale help

OpenStudy (mathmale):

@Hotchellerae21: I strongly encourage you to put to use what you learned from the previous problem to solve this present one. "4 yards per carry" is a rate:

OpenStudy (mathmale):

\[\frac{ 4~yards }{ 1~carry }\]

OpenStudy (anonymous):

i don't need help anymore i got it thanks

OpenStudy (mathmale):

Also, you are told that this guy carried the ball 16 times. Note that your answer MUST have the unit of measurement "yards."